He is the question below. I need to define the variable for x and y, write the appropriate equations, and solve the equation by either substitution, addition, or graphically. THX!
- The difference of two numbers is nine more than the smaller number, and the sum of the two numbers is 54. Find the smaller number.

Answer by xcentaur(357)

let the larger number be x
let the smaller number be y


difference of two numbers means (x-y)
this difference is nine more than the smaller number
x-y=9+y
x-y-y=9
x-2y=9......[1]


the sum of the two numbers is 54
x+y=54.......[2]


take the two equations together,
x-2y=9
x+y=54
subtract equation [2] from [1]
x-2y-(x+y)=9-54
x-2y-x-y=-45
-3y=-45
3y=45
y=15


x+y=54
x=54-y=54-15=39


x=39,y=15


Hope this helps,
Prabhat